OpenStudy (anonymous):

Marcus purchased 80 shares of stock in a computer company at $74.19 per share and Taylor purchased 65 shares of stock in a different computer company for $85.21 per share. After holding the stock for two years, Marcus sold his for a total of $6,404.57, and Taylor sold hers for a total of $6,192.83. Which person had a higher ROI and by how much? Marcus: 3.9% more Taylor: 3.9% more Marcus: 7.9% more Taylor: 11.8% more

OpenStudy (anonymous):

@NaCl

OpenStudy (anonymous):

@ivettef365 @mathmate

OpenStudy (mathmate):

ROI=(selling price - purchase price)/purchase price Example: for Marcus, ROI=(( 6,404.57/80)-74.19)/74.19=7.91%

OpenStudy (anonymous):

so now what??

OpenStudy (mathmate):

Calculate the ROI for Taylor and compare which is more profitable.

OpenStudy (anonymous):

I got 0.983 for Taylor is that correct @mathmate

OpenStudy (mathmate):

Not quite. What is the expression to get that?

OpenStudy (anonymous):

6129.83-5538.65/6129.83 = 0.096 I tried it again and got this

OpenStudy (mathmate):

Taylor's investment is 5538.65, so you really want to divide by his investment, instead of the market value.

OpenStudy (anonymous):

i dont understand im sorry

OpenStudy (mathmate):

The definition of ROI is gain/cost gain is 6192.83-5538.65 (as you had it) but cost is what he paid, which is 5538.65 so the ROI = (6192.83-5538.65)/5538.65

OpenStudy (anonymous):

Ok gotcha, I got 0.118

OpenStudy (mathmate):

Now you only have to compare the two ROI and see which one is more profitable.

OpenStudy (anonymous):

So how would I do that

OpenStudy (mathmate):

The question asks who has a higher ROI, can you tell me who has a higher ROI?

OpenStudy (mathmate):

Note: ROI means return on investment.

OpenStudy (mathmate):

And after that, you would subtract the lower from the higher ROI and then make your choice from the 4 choices given.

OpenStudy (anonymous):

Oh Taylor ? So it would be D ?

OpenStudy (mathmate):

Ouch! They did not ask what the ROI's are. They are asking the difference between the two ROIs.

OpenStudy (anonymous):

ITS C. 7.9

OpenStudy (mathmate):

7.9 is Marcus's ROI, not the difference between the 2. You have to choose from the remaining choices depending on whose ROI is higher. Go ahead!

OpenStudy (anonymous):

i want to say B

OpenStudy (mathmate):

because...

OpenStudy (anonymous):

Because Taylors ROI is higher?

OpenStudy (mathmate):

Right! Good job!

OpenStudy (anonymous):

thanks so much for your help i appreciate it!

OpenStudy (mathmate):

You're very welcome! Wish you great success with your studies!